### v1

# ALTo: Adaptive-Length Tokenizer for Autoregressive Mask Generation

Abstract
While humans effortlessly draw visual objects and shapes by adaptively allocating attention based on their complexity, existing multimodal large language models (MLLMs) remain constrained by rigid token representations. Bridging this gap, we propose ALTo, an adaptive length tokenizer for autoregressive mask generation. To achieve this, a novel token length predictor is designed, along with a length regularization term and a differentiable token chunking strategy. We further build ALToLLM that seamlessly integrates ALTo into MLLM. Preferences on the trade-offs between mask quality and efficiency is implemented by group relative policy optimization (GRPO). Experiments demonstrate that ALToLLM achieves state-of-the-art performance with adaptive token cost on popular segmentation benchmarks. Code and models are released at https://github.com/yayafengzi/ALToLLM.
Figure 1: ALToLLM realizes adaptive-length mask token generation according to object complexity.
1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in image and text understanding tasks. However, their generative abilities remain largely limited to text liu2023llava ; chen2024internvl ; bai2023qwenvl ; yuan2024minicpmv ; wu2024deepseek ; grattafiori2024llama . Given the inherent differences between text and image modalities, introducing a lightweight image decoder into multimodal understanding models to enable image generation remains a significant challenge. To align with the next-token prediction paradigm of text generation, discretizing images using image tokenizers (e.g., VQGAN esser2020taming ) has become a natural and effective approach. Within this framework, various visual modalities—including RGB images, segmentation masks, and depth maps—can be uniformly represented as “images”, enabling unified modeling for both multimodal understanding and generation lu2023unifiedio ; Lu2023unifiedio2 .
Early image tokenizers typically represent images using fixed-length token sequences without considering the inherent complexity of the images esser2020taming ; van2017vqvae ; yu2024titok ; ge2023planting . This fixed-length design may lead to insufficient representation for complex images while generating redundant tokens for simpler ones, resulting in resource wastage and reduced efficiency. In contrast, humans can flexibly allocate attention based on the complexity of the task legg2007universal . For example, segmenting complex shapes requires more attention and effort compared to simpler shapes.
Recent arts are dedicated to learning hierarchical and flexible tokens roman2025flextok ; Wang2024Emu3 ; wang2025himtok . The representations become increasingly fine-grained as the number of tokens increases. Based on our observations, the number of tokens required to represent fine-grained edge shapes can vary drastically depending on their complexity.
Table 1: The flexibility and autonomous adaptivity of different token representation methods. Flexibility refers to hierarchical coarse-to-fine token representation, while autonomous adaptivity denotes spontaneous allocation of token numbers based on object complexity.
Token Representation Flexibility Autonomous Adaptivity
VAE kingma2013vae $\times$ $\times$
VQVAE van2017vqvae $\times$ $\times$
VQGAN esser2020taming $\times$ $\times$
TiTok yu2024titok $\times$ $\times$
FlexTok roman2025flextok $\checkmark$ $\times$
Emu3 Wang2024Emu3 $\checkmark$ $\times$
Chameleon Lu2023Chameleon $\times$ $\times$
HiMTok wang2025himtok $\checkmark$ $\times$
ALIT duggal2025adaptive $\times$ $\times$
ElasticTok yan2025elastictok $\checkmark$ $\times$
ALToLLM (ours) $\checkmark$ $\checkmark$
In recent years, several studies duggal2025adaptive ; yan2025elastictok have explored adaptive-length tokenization for image representations. The problem, however, is that they all determine adaptive lengths by relying on heuristic rules conditioned on the input image, rather than allowing the model to decide on its own. Although this is feasible in image tokenization, it becomes impractical for image generation since the reconstruction loss is unavailable. As a result, it becomes imperative to enable the model to autonomously determine the adaptive token length, specifically for MLLM scenarios that are expensive in computation, like MLLM-based object segmentation.
To enable adaptive-length modeling for the specific task of mask image generation, we propose ALTo, an Adaptive-Length Tokenizer designed for autoregressive mask generation. We further develop ALToLLM. As shown in Fig. 1, ALToLLM is a multimodal large language model (MLLM) that realizes instruction-based mask generation using adaptive-length mask tokens according to object complexity. At the core of ALTo is a novel Token Length Predictor (TLP) embedded within an encoder–VQ–decoder architecture. Given an input mask image, the ALTo encoder is responsible not only for generating discrete tokens but also for predicting the appropriate token sequence length via TLP. To support adaptive-length learning, we introduce a length regularization term and a differentiable token chunking strategy. Together, these enable ALTo to effectively encode masks into variable-length token sequences. To evaluate the effectiveness of ALTo, we construct ALToLLM without any bells and whistles, making no modifications to the underlying LLM architecture or training paradigm. The model is trained using supervised fine-tuning and Group Relative Policy Optimization (GRPO) on referring image segmentation tasks. ALToLLM learns to adaptively insert an end-of-mask token (<ALTo_End>) once sufficient mask tokens have been generated. Moreover, GRPO allows dynamic control over token length to balance mask quality and computational efficiency.
In summary, the contributions are as follows:
  • We propose ALTo, an adaptive-length mask tokenizer that, for the first time, enables the model to autonomously determine the number of mask tokens based on the complexity of the input mask.
  • We develop ALToLLM, which integrates ALTo into a multimodal large language model (MLLM), enabling adaptive mask token generation for object segmentation tasks. The number of generated tokens can vary from as few as 2 to as many as 32, with most cases around 16, allowing ALToLLM to balance quality and efficiency under different scenarios via GRPO.
  • Extensive experiments demonstrate that ALTo enables effective and efficient mask image reconstruction, while ALToLLM achieves state-of-the-art performance with adaptive token usage across various object segmentation benchmarks, including referring expression segmentation and open-vocabulary segmentation.
2 Related work
Visual tokenzers play important roles in various visual tasks, such as image reconstruction van2017vqvae ; esser2020taming , visual compression yan2025elastictok , and visual generation ramesh2021zero ; tian2024var ; lu2023unifiedio ; wang2025himtok . VQVAE van2017vqvae ; razavi2019generating and VQGANesser2020taming are popular frameworks that encode images into discrete 2D tokens by vector quantization. BEiT bao2021beit exploits visual tokens in masked image modeling. To reduce the redundancy in 2D space, TiTok yu2024titok and SEED ge2023planting produce 1D sequence for image tokenization. Methods above typically use a rigid number of tokens to represent images, regardless of the complexity of the visual content. To have flexible tokenization, FlexTok roman2025flextok projects images into 1D variable-length token sequences. ElasticTok yan2025elastictok proposed an adaptive tokenizer for images and videos by dropping a random number of latter tokens during training. ALIT duggal2025adaptive discretizes images into flexible-length tokens by recurrent distillation until the reconstruction quality is okay or meet the maximum iterations. HiMTok wang2025himtok learns 1D hierarchical mask tokens to represent segmentation masks from coarse to fine. However, these methods cannot decide an adaptive number of tokens autonomously. Trials have been made by heuristic rules about image reconstruction quality duggal2025adaptive ; yan2025elastictok , which increases computational overhead and becomes impossible for image generation tasks. Our proposed ALTo is both flexible and adaptive, as illustrated in Tab.  1.
MLLM-based image segmentation. Recent MLLM-based segmentation methods primarily follow three paradigms lan2024text4seg ; wang2025himtok ; tang2025ufo . MLLM-segmentation joint models such as LISA lai2024lisa , GSVA xia2024gsva , GLaMM rasheed2024glamm , PixelLM ren2024pixellm , and PSALM zhang2024psalm , create semantic-to-pixel connections through LLM hidden states and rely on additional segmentation modules. Text-based methods, including Text4Seg lan2024text4seg , LLaFS zhu2024llafs and VistaLLM pramanick2024jack , represent masks as text sequences (pixel classes or polygon vertices), suffering from heuristic and inaccurate mask representation. Interestingly, segmentation masks could also be viewed as images so that we can rethink image segmentation as a mask generation task lu2023unifiedio ; Lu2023unifiedio2 ; bar2022visual . HiMTok wang2025himtok applies the idea by utilizing a hierarchical mask tokenizer into LLMs. Going a step further, ALToLLM generates adaptive-length token sequences, which is efficient and effective.
Reinforcement learning. Reinforcement learning (RL) has become increasingly important for enhancing vision-language models setlur2024rl ; ma2024coevolving ; qu2025latent ; li2024getting ; zhai2025enhancing . Approaches like direct preference optimization rafailov2023direct and proximal policy optimization john2017ppo face challenges with data efficiency and reward stability. Group relative policy optimization (GRPO) shao2024deepseekmath has emerged as a promising alternative through its groupwise reward mechanism. Recent applications demonstrate GRPO’s effectiveness across various vision-language tasks. Visual-RFT liu2025visualrft combines GRPO with verifiable rewards for efficient model adaptation. Vision-R1 huang2025visionr1 employs GRPO with progressive thinking suppression for complex reasoning. Seg-Zero liu2025segzero achieves zero-shot segmentation through pure RL. These works apply RL to text output, while we make it for preference optimization on mask token output.
3 Methods
3.1 Overview
Figure 2: Training recipes for ALTo and ALToLLM. (a) ALTo Pretraining: Joint training of mask tokenizer (MT) and de-tokenizer (MD); (b) Adaptive-length Prediction: Training only the token length predictor (TLP); (c) Multimodal Integration: Exclusive training of MLLM with frozen ALTo for language-aware adaptation; (d) Group Relative Policy Optimization: Reinforcement learning for MLLM optimization. Input image in (b), (c) and (d) is processed identically to (a), omitted for visual clarity.
The proposed adaptive length tokenizer (ALTo) represents object masks as token sequences whose lengths adapt to the complexity of objects autonomously. Simple objects (e.g., a sphere) may require few tokens, while intricate structures (e.g., complicated shapes and multiple objects) may use up to 32 tokens. Built on this, ALToLLM is introduced to perform instruction mask generation for referring image segmentation, as shown in Fig. 3. We design a multi-stage training recipe for ALTo and ALToLLM to learn flexible, adaptive, and effective mask representations and achieve strong segmentation performance, as shown in Fig. 2.
Inference. As illustrated in Fig. 3, ALToLLM takes as input the image and text by the popular ViT-projector-LLM architecture liu2023llava ; chen2024far , then autoregressively generates both text tokens and compact mask tokens of adaptive length. The mask tokens along with the pixel-encoded features are fed into the mask de-tokenizer to generate the final mask.
Training. As shown in Fig. 2, the training recipe consists of three progressive stages. Stage 1: We pretrain the mask tokenizer (MT) and mask de-tokenizer (MD) to reconstruct complex masks using variable-length tokens. Stage 1.5: We fine-tune the token length predictor (TLP) to enable adaptive tokenization. Stage 2: We leverage ALTo to generate both fixed-length and adaptive-length token labels, which are used to supervise ALToLLM. This equips the model with the basic ability to understand and generate both text and mask tokens. Stage 3: We employ GRPO shao2024deepseekmath to further adjust specific preferences on trade-off between mask quality and token efficiency. We will introduce the details in the following subsections.
3.2 ALTo
The Adaptive-Length Tokenizer (ALTo) comprises three components: a mask tokenizer (MT), a mask de-tokenizer (MD) with a pixel encoder, and a token length predictor (TLP), as shown in Fig. 2 (a). Following HiMTok wang2025himtok , MT utilizes a transformer encoder with 32 learnable latent tokens to extract information from the input mask and then discretized into 32 mask tokens via vector quantizer(VQ). To support variable-length tokenization, a random number of tail tokens are dropped during training, retaining only the leading tokens. MD is a bidirectional transformer. Differently from HiMTok, MD takes as input the mask tokens and 256 pixel-encoded image features rather than learnable latent tokens. This provides fine-grained guidance for mask generation, inspired by UViM kolesnikov2022uvim . In training stage 1, the reconstruction is supervised by a mean squared error (MSE) loss $\mathcal{L}_{\text{Mask}}$ to pretrain MT and MD.
The novel token length predictor (TLP) determines the optimal number of tokens for each mask. TLP leverages the CLS token feature $T_{\text{cls}}$, which encodes global image features, together with the 32 mask token features $T\in\mathbb{R}^{32\times d}$, to predict a proper token length. $T_{\text{cls}}$ is used as a query in an attention mechanism to evaluate the importance of each mask token. For each mask token $T_{i}$, a gated key is generated by SwiGLU as $k_{i}=(W_{\text{v}}T_{i})\odot\sigma(W_{\text{g}}T_{i})$. The probability for each token to be the stopping point is computed via scaled dot-product attention: $p=\text{softmax}(q_{\text{cls}}k^{T}/\sqrt{d})$. The predicted length is computed as the mathematical expectation $\hat{L}=\sum_{i=1}^{32}i\cdot p_{i}$.
Accordingly, the first $\hat{L}$ tokens are selected and sent to the MD, while the remaining tokens are zero-padded, represented as $H\bigodot T$, where $H$ is a binary mask defined as $H=\mathbb{I}[i\leq\hat{L}]$. However, such token chunking strategy is not differentiable, which prevents gradients from the mask de-tokenizer from flowing back to the TLP. To address this, we introduce a differentiable token chunking strategy by considering the stopping probability distribution $p$. The probability that the $i$-th token is used is given by the cumulative probability $P_{i}=\sum_{j\geq i}p_{j}$, which indicates that the stop position is later than this token and provides a soft version of token chunking. This inspires us to apply a straight-through estimator as $\hat{T}=(P-\hat{P}+H)\bigodot T$. In this formulation, the predicted mask is then given by $M_{\text{pred}}=\text{MD}(\hat{T},X_{\text{img}})$.
In stage 1.5, we use a reconstruction loss $\mathcal{L}_{\text{Mask}}=\text{MSE}(M_{\text{pred}},M_{\text{gt}})$ to optimize mask reconstruction, and a length regularization term $\mathcal{L}_{\text{Length}}=-\lambda\hat{L}$ to encourage shorter token sequences, balancing accuracy and efficiency while MT and MD are frozen. The final combined loss is $\mathcal{L}_{\text{ALA}}=\mathcal{L}_{\text{Mask}}+\mathcal{L}_{\text{Length}}$. Further details about the length supervision design are provided in the Appendix. A.
Figure 3: Architecture of the proposed ALToLLM.
3.3 ALToLLM
ALToLLM is built naturally on MLLM architecture, and learned by supervised fine-tuning (SFT) and group relative policy optimization (GRPO).
During SFT (stage 2), ALToLLM receives adaptive-length mask tokens provided by the frozen ALTo module, along with text and image tokens. ALToLLM is supervised using two objectives: a cross-entropy loss $\mathcal{L}_{\text{ce}}$ for next-token prediction across the multimodal sequence, and a mask prediction accuracy loss $\mathcal{L}_{\text{mask}}$, which combines binary cross-entropy loss and Dice loss to ensure precise mask reconstruction. This dual-objective training enables ALToLLM to effectively align textual and visual information, and to autoregressively generate both language and adaptive-length mask tokens, which shows the effectiveness of ALTo.
We employ GRPO in stage 3 to adjust trade-off preferences flexibly based on the model after stage 2.
1) Group sampling: For each input consisting of an image and text, we sample $g$ multimodal responses from ALToLLM. A valid $i$-th sample must contain the following token sequence, where $L_{i}$ denotes the adaptive length:
[…]
2) Reward computation: The composite reward $R_{i}$ for the $i$-th sample consists of three components:
$R_{i}$ $=\underbrace{\mathbb{I}_{\text{format}}}_{R_{\text{valid}}}+\underbrace{\text{IoU}}_{R_{\text{accuracy}}}-\underbrace{\alpha L_{i}}_{R_{\text{efficiency}}},$ (2)
where $R_{\text{valid}}$ is $1$ if the sample strictly follows the format in Eq. 1, and 00 otherwise. $R_{\text{accuracy}}$ is the intersection-over-union (IoU) score between the predicted and ground truth masks, with the predicted mask reconstructed by the MD using the adaptive mask tokens. $R_{\text{efficiency}}$ is a linear penalty $-\alpha L_{i}$ proportional to the token length $L_{i}\in\{1,\ldots,32\}$, scaled by the trade-off parameter $\alpha\geq 0$.
3) Relative policy optimization: We optimize the policy using the clipped objective $\mathcal{J}_{\text{GRPO}}(\theta)$:
$\mathbb{E}\left[\frac{1}{g}\sum_{i=1}^{g}\min\left(\frac{\pi_{\theta}}{\pi_{\text{old}}}A_{i},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\text{old}}},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\right],$ (3)
where $\pi_{\theta}$ is the current policy being optimized (parameterized by $\theta$), $\pi_{\text{old}}$ is the policy before the update (used for importance sampling), $\pi_{\text{ref}}$ is the reference policy (typically the initial supervised policy), $A_{i}=(R_{i}-R_{\text{mean}})/R_{\text{std}}$ is the normalized advantage computed within each group, $D_{\text{KL}}$ is the Kullback-Leibler (KL) divergence enforcing policy stability, $\epsilon$ is the clip range (usually 0.1-0.3) controlling update aggressiveness, and $\beta$ is the KL penalty coefficient balancing exploration and constraint.
4 Experiments
4.1 Experimental settings
Datasets. For stages 1 and 1.5, we construct the training and validation sets of Multi-Target-SA1B from the SA1B dataset by randomly selecting multiple masks from all annotations for each image. Examples from Multi-Target-SA1B are shown in Fig. 4. This approach yields complex multi-target masks, facilitating the learning of expressive mask representations by ALTo. For stage 2, we used all HiMTok and Multi-Target-SA1B datasets for SFT. For Multi-Target-SA1B, we input the bounding boxes of all targets as “<box>[[],[],...]</box>”. To ensure that the model supports both fixed-length and adaptive-length prompts, we randomly assign half of the data to each prompt type, as detailed in the Appendix. B. For stage 3, we use Multi-Target-SA1B, the RefCOCO series yu2016modeling ; mao2016generation , and gRefCOCO liu2023gres to maintain complex mask representation and language understanding during RL.
Figure 4: Examples from the Multi-Target-SA1B dataset.
Table 2: Performance comparison on gRefCOCO. We report cIoU, gIoU and average token length. FT indicates fine-tuning on referring expression data.
Method val testA testB
cIoU gIoU Length cIoU gIoU Length cIoU gIoU Length
LISA-7B lai2024lisa 38.7 32.2 - 52.6 48.5 - 44.8 39.7 -
LISA-7B (FT) lai2024lisa 61.8 61.6 - 68.5 66.3 - 60.6 58.8 -
GSVA-7B xia2024gsva 61.7 63.3 - 69.2 70.1 - 60.3 61.3 -
GSVA-7B (FT) xia2024gsva 63.3 66.5 - 69.9 71.1 - 60.5 62.2 -
GroundHog-7B zhang2024groundhog - 66.7 - - - - - - -
SAM4MLLM-8B chen2024sam4mllm 67.8 71.9 - 72.2 74.2 - 63.4 65.3 -
UniRES++ 2025liuunires++ 69.9 74.4 - 74.5 76.0 - 66.6 69.8 -
[…]
ALToLLM-8B (FL) 74.8 77.6 32 78.5 78.7 32 76.4 76.7 32
ALToLLM-8B (AL) 75.4 78.0 17.5 78.8 78.9 19.4 76.6 76.9 17.3
Implementation details. ALTo processes input and reconstructs masks at $256\times 256$ resolution. During training and inference, the MLLM processes images at $448\times 448$, while the pixel encoder encodes image at $1024\times 1024$. In stage 1, MT and MD are initialized from TiTok-L-32 yu2024titok , and the pixel encoder is initialized from SAM-ViT-L kirillov2023sam . In stage 1.5, the feature dimension of TLP is set to 1024, consistent with MT. The length penalty coefficient is set to 0.0001, 0.001, 0.01, or 0.1, among which 0.01 is found to be optimal in subsequent experiments and is chosen for later stages. In stage 2, ALToLLM-8B is initialized from InternVL-2.5-8B chen2024far . Stage 3 trains the RL model based on the stage 2 checkpoint, with the length penalty set to 1e-2, 5e-3, 3e-3, 2e-3, 1e-3, or 1e-4, which are compared in later experiments. The KL penalty is set to 1e-3. We sample 12 group responses with a temperature of 1 and top-k of 10. Stages 1, 1.5, and 3 are trained on 8$\times$×A100 GPUs (80GB each), and stage 2 on 16$\times$×A100 GPUs. Training durations are: stage 1 for 2 days, stage 1.5 for 3 hours, stage 2 for 5 days, and stage 3 for 5 hours.
4.2 Comparative results
We evaluate ALToLLM-8B on the following tasks, considering two variants: (1) a fixed-length version (ALToLLM-8B (FL)) using 32 mask tokens, and (2) an adaptive-length version (ALToLLM-8B (AL)) that dynamically generates 1–32 mask tokens.
Generalized referring expression segmentation with multiple targets. We evaluate ALToLLM-8B on the generalized referring expression segmentation task (gRefCOCO liu2023gres ) and achieve state-of-the-art performance, as shown in Table 2. This task requires language-guided segmentation with multi-target referring expressions, demonstrating our model’s ability to learn complex mask representations. Compared to the fixed-length variant (ALToLLM-8B (FL)), the adaptive-length variant (ALToLLM-8B (AL)) achieves higher performance and significantly reduces average token length, indicating improved segmentation accuracy and token efficiency. The superior performance of the adaptive-length approach may be attributed to its ability to use only the necessary tokens for simple masks, avoiding the noise introduced by redundant tokens. We also compare the average generation time per sample of the two variants. As shown in Table 3, the adaptive length variant achieves a consistently shorter generation time in all splits.
Table 3: Comparison of average generation time per sample (in seconds) between fixed-length and adaptive-length variants on gRefCOCO. Generation time is measured on a single A100 GPU with batch size 1.
Method val testA testB
ALToLLM-8B (FL) 1.079 1.079 1.076
ALToLLM-8B (AL) 0.710 0.753 0.669
Referring expression segmentation. We further evaluate ALToLLM-8B on referring expression segmentation, a single-target version of the generalized task. Experiments are conducted on three standard benchmarks: RefCOCO yu2016modeling , RefCOCO+ yu2016modeling , and RefCOCOg mao2016generation . As shown in Table 4, ALToLLM-8B achieves state-of-the-art results across all datasets.
Figure 5: Examples from the constructed multi-class version of open-vocabulary segmentation datasets. (a) ADE20K (A-150); (b) PASCAL Context59 (PC-59); (c) PASCAL VOC 20 (PAS-20).
Table 4: Performance comparison on RefCOCO, RefCOCO+, and RefCOCOg. We report cIoU. FT indicates fine-tuning on referring expression data.
Methods RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val (U) test (U)
Text4Seg-InternVL2-8B lan2024text4seg 79.2 81.7 75.6 72.8 77.9 66.5 74.0 75.3
PolyFormer - B liu2023polyformer 74.8 76.6 71.1 67.6 72.9 59.3 67.8 69.1
VistaLLM - 7B pramanick2024jack 74.5 76.0 72.7 69.1 73.7 64.0 69.0 70.9
LISA - 7B lai2024lisa 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.4
LISA - 7B (FT) lai2024lisa 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
PixelLM - 7B ren2024pixellm 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5
GSVA - 7B xia2024gsva 76.4 77.4 72.8 64.5 67.7 58.6 71.1 72.0
GSVA - 7B (FT)  xia2024gsva 77.2 78.9 73.5 65.9 69.6 59.8 72.7 73.3
PSALM zhang2024psalm 83.6 84.7 81.6 72.9 75.5 70.1 73.8 74.4
GLaMM rasheed2024glamm 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9
GroundHog - 7B zhang2024groundhog 78.5 79.9 75.7 70.5 75.0 64.9 74.1 74.6
SAM4MLLM - 8B chen2024sam4mllm 79.8 82.7 74.7 74.6 80.0 67.2 75.5 76.4
M²SA - 7B jang2025mmr 74.0 76.8 69.7 63.1 67.2 56.1 67.0 68.3
AnyRef he2024anyref 74.1 75.5 70.8 64.1 68.7 57.5 68.1 69.9
AnyRef (FT) he2024anyref 76.9 79.9 74.2 70.3 73.5 61.8 70.0 70.7
SegAgent - LLaVA+SAM zhu2025segagent 79.2 81.4 75.7 71.5 76.7 65.4 74.8 74.9
SegAgent - Qwen+SAM zhu2025segagent 78.0 80.3 75.0 70.9 75.5 65.8 74.5 74.6
SegAgent - LLaVA+SClick zhu2025segagent 77.8 80.0 74.1 66.7 71.2 59.9 70.5 71.3
SegAgent - Qwen+SClick zhu2025segagent 79.7 81.4 76.6 72.5 75.8 66.9 75.1 75.2
Seg-Zero-7B liu2025segzero - 80.3 - - 76.2 - - 73.6
[…]
ALToLLM-8B (FL) 84.9 85.4 83.9 81.4 83.8 77.9 80.4 80.7
ALToLLM-8B (AL) 85.8 86.6 84.7 81.3 83.8 77.0 80.6 81.4
Multi-granularity segmentation. We evaluate ALToLLM-8B on RefCOCOm 2025liuunires++ , a multi-granularity referring segmentation dataset containing both part-level and object-level referring expressions. As shown in Table 5, ALToLLM-8B (AL) achieves the best performance.
Table 5: Performance comparison on RefCOCOm. We report mIoU for part-level and object & part-level expressions. indicates results reproduced by us using the official code and settings.
Methods val testA testB
Part Obj & Part Part Obj & Part Part Obj & Part
X-Decoder zou2023xdecoder 16.2 29.5 13.6 23.6 20.3 33.8
SEEM zou2023seem 16.1 29.4 13.6 23.4 20.4 33.9
UniRES wang2024unires 19.6 34.3 16.4 27.8 25.2 41.7
LISA-7B lai2024lisa 21.3 34.3 18.5 28.6 25.7 40.1
GSVA-7B xia2024gsva 11.4 23.1 9.2 19.2 16.8 28.2
GLaMM rasheed2024glamm 21.4 35.3 18.6 29.5 26.9 41.1
M²SA-7B jang2025mmr 22.4 35.5 19.9 30.1 27.1 41.4
[…]
ALToLLM-8B (FL) 25.5 39.2 22.6 33.3 30.3 46.7
ALToLLM-8B (AL) 25.5 39.1 22.6 33.2 30.2 46.5
Multi-class open-vocabulary segmentation. To demonstrate our model’s ability to segment multiple and complex targets in open-vocabulary scenarios, we construct a multi-class version of open-vocabulary segmentation datasets by randomly merging annotations from several classes, including ADE20K (A-150) zhou2019semantic , PASCAL Context59 (PC-59) mottaghi2014role , and PASCAL VOC 20 (PAS-20) everingham2010pascal , as shown in Fig. 5. We reproduce the inference pipelines for LISA lai2024lisa and M²SA jang2025mmr for comparison. As shown in Table 7, ALToLLM-8B achieves state-of-the-art results.
Table 6: Performance comparison on multi-class open-vocabulary segmentation. We report gIoU.
Method A-150 PC-59 PAS-20
LISA-7B lai2024lisa 39.1 53.3 67.8
M²SA-7B jang2025mmr 63.9 74.1 79.5
ALToLLM-8B (FL) 65.6 75.2 81.2
ALToLLM-8B (AL) 65.8 75.4 81.1
Table 7: Ablation study of the pixel encoder on the Multi-Target-SA1B validation dataset. We report gIoU.
Method Pixel Encoder gIoU
ALTo (Ours) 94.4
ALTo (Ours) $\times$ 91.9
4.3 Adaptive length preference adjustment via reinforcement learning
Figure 6: Metrics of sampled responses during GRPO training. (a) Average token length, (b) Average IoU, (c) Generation entropy.
Figure 7: Comparison of token cost between fixed-length and adaptive-length models with different length preferences. FL denotes the fixed-length model from stage 2. AL denotes stage 3 models trained with different length penalties (from left to right: 1e-2, 5e-3, 3e-3, 2e-3, 1e-3, 1e-4). (a) Multi-Target-SA1B val, (b) gRefCOCO val, (c) Multi-Class A-150 (zero-shot).
Figure 8: Analysis of the penalty coefficient in stage 1.5. (a) Length metrics for different penalty coefficients. (b) Distribution of length metrics for penalty coefficients 0.0001, 0.001, 0.01, and 0.1.
By tuning the length penalty in the reward function, we can flexibly control the model’s preference for adaptive token lengths while maintaining high IoU within a few hundred training steps by GRPO. As the average adaptive length decreases, the generation entropy also decreases, indicating that later tokens are associated with higher uncertainty. This trend is visualized in Fig. 6. To quantify the token savings achieved by adaptive-length tokens at various IoU levels, we compare six models trained with different length penalties in stage 3, alongside a fixed-length baseline from stage 2. Validation is conducted on Multi-Target-SA1B for complex mask representation and gRefCOCO liu2023gres for language understanding. For zero-shot evaluation, we use the multi-class A-150 dataset, which is not seen during stage 3 training. As shown in Fig. 7, adaptive-length models consistently save more than 10 tokens at the same IoU level, and this advantage persists even in zero-shot scenarios.
4.4 Ablation study on ALTo
We demonstrate that the pixel encoder is necessary for the reconstruction of complex masks. As shown in Table 7, removing the pixel encoder leads to a significant drop in reconstruction gIoU on the Multi-Target-SA1B validation dataset, indicating that pixel-level features are crucial for capturing fine-grained mask details. We compare different penalty coefficients during stage 1.5 training and analyze their impact on adaptive-length prediction to select an appropriate value, as shown in Fig. 8. The results indicate that a coefficient of 0.01 achieves a good balance between IoU and a wide range of predicted lengths, with high standard deviation and entropy. Therefore, we adopt this value for stage 2 training.
5 Conclusions
We present ALToLLM, an innovative framework that dynamically adapts the number of mask tokens according to object complexity. ALToLLM approaches mask tokens as a visual language system, where our ALTo intelligently determines the optimal token count for each object. Furthermore, by integrating ALTo with MLLMs, the system can interpret linguistic descriptions and correspondingly adjust token allocation. Extensive experiments demonstrate state-of-the-art performance on RefCOCO yu2016modeling , RefCOCO+ yu2016modeling , RefCOCOg mao2016generation , RefCOCOm 2025liuunires++ , and gRefCOCO liu2023gres benchmarks, validating our approach’s effectiveness in aligning linguistic expressions with adaptive mask tokenization.
Appendix
Appendix A Details of TLP
The Adaptive-Length Tokenizer (ALTo) utilizes a token length predictor (TLP) to dynamically determine the number of tokens required for each mask. In this section, we provide further mathematical details and training insights to clarify the differentiable token chunking strategy.
The overall loss function consists of a reconstruction loss and a length regularization term:
$\mathcal{L}_{\text{ALA}}=\mathcal{L}_{\text{Mask}}+\mathcal{L}_{\text{Length}},$
where $\mathcal{L}_{\text{Length}}=-\lambda\hat{L}$ penalizes longer token sequences. The gradient of the loss is given by
$\nabla\mathcal{L}_{\text{ALA}}=$ $\nabla\mathcal{L}_{\text{mask}}+\nabla\mathcal{L}_{\text{length}}$
$=$ $\sum_{i}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{i}}\nabla\hat{T}_{i}+\lambda\sum_{i}i\nabla p_{i}$
If we directly chunk the token sequence, such as $\hat{T}=H\bigodot T$, we obtain $\nabla\hat{T}=0$ because $H$ is non-differentiable and $T$ is generated by a frozen mask tokenizer. As a result, the gradient of $\mathcal{L}_{\text{mask}}$ cannot be backpropagated to the TLP, and only the token length is optimized to be minimal.
To address this limitation, we introduce a differentiable token chunking strategy. Specifically, we use the cumulative probability $P_{i}=\sum_{j\geq i}p_{j}$ to represent the probability that the $i$-th token is used, and $\hat{P}$ denotes the detached version of $P$. We then construct a soft token chunking as $\hat{T}=(P-\hat{P}+H)\bigodot T$. In this way, we have $\nabla\hat{T}=\nabla P\bigodot T$. The overall gradient is given by
$\nabla\mathcal{L}_{\text{ALA}}=$ $\nabla\mathcal{L}_{\text{mask}}+\nabla\mathcal{L}_{\text{length}}$
$=$ $\sum_{i}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{i}}P_{i}T_{i}+\lambda\sum_{i}i\nabla p_{i}$
$=$ $\sum_{i}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{i}}T_{i}\sum_{j\geq i}\nabla p_{j}+\lambda\sum_{i}i\nabla p_{i}$
$=$ $\sum_{i}\nabla p_{i}\sum_{j\leq i}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda\sum_{i}i\nabla p_{i}$
$=$ $\sum_{i}\nabla p_{i}(\sum_{j\leq i}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda i)$
for the best predict length $\hat{L}=k$, there is $(\sum_{j\leq k}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda k)<(\sum_{j\leq k-1}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda{(k-1)})$ and $(\sum_{j\leq k}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda k)<(\sum_{j\leq k+1}\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{j}}T_{j}+\lambda{(k+1)})$ , which means
$-\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k+1}}T_{k+1}<\lambda<-\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k}}T_{k}$
This form shows that the model is encouraged to select the minimal number of tokens that still achieve high reconstruction quality, as the regularization term $\lambda$ acts as a threshold for including additional tokens.
Intuitively, the model will only increase the predicted token length if the marginal gain in reconstruction quality outweighs the regularization penalty. This mechanism is analogous to a reward-cost trade-off in reinforcement learning, where the "reward" for including an additional token must exceed the cost $\lambda$.
Appendix B Prompt design
We prepared different prompt templates for instruction tuning on adaptive-length and fixed-length segmentation. Tabs 9 and 9 are the templates for adaptive-length ALToLLM tuning, and Tabs 11 and 11 are the templates for fixed-length version.
Table 8: Templates of instruction for adaptive-length segmentation.
“Segment <ref>{}</ref> by adaptive length." “Create a mask for <ref>{}</ref> by adaptive length." “Generate a mask for <ref>{}</ref> by adaptive length." “Do segmentation for <ref>{}</ref> by adaptive length." “Please give the mask for <ref>{}</ref> by adaptive length." “What is the mask for <ref>{}</ref> by adaptive length?" “Can you segment <ref>{}</ref> by adaptive length?"
Table 9: Templates of response for adaptive-length segmentation.
“The adaptive mask appears at {}." “The adaptive mask is created as {}." “I can generate the adaptive mask at {}." “The adaptive mask is {}." “I can give the adaptive mask at {}." “Its adaptive mask located at {}." “Sure, the adaptive mask is {}."
Table 10: Templates of instruction for fixed-length segmentation.
“Segment <ref>{}</ref>"̇ “Create a mask for <ref>{}</ref>"̇ “Generate a mask for <ref>{}</ref>"̇ “Do segmentation for <ref>{}</ref>"̇ “Please give the mask for <ref>{}</ref>"̇ “What is the mask for <ref>{}</ref>?" “Can you segment <ref>{}</ref>?"
Table 11: Templates of response for fixed-length segmentation.
“The mask appears at {}." “The mask is created as {}." “I can generate the mask at {}." “The mask is {}." “I can give the mask at {}." “Its mask located at {}." “Sure, the mask is {}."
Appendix C Application examples
Fig. 9 illustrates examples of ALToLLM in practical applications. As demonstrated, ALToLLM effectively segments the target object referred to by the user by leveraging adaptive-length mask tokens. This approach not only ensures high segmentation accuracy but also optimizes computational efficiency by dynamically adjusting the number of tokens based on object complexity.
Figure 9: Examples of ALToLLM with adaptive-length segmentation.